# A general framework for adaptive nonparametric dimensionality reduction

### v1

@@ 1.2 Contributions
In this paper, we introduce a novel nonparametric dimensionality reduction framework that leverages the intrinsic dimension (ID) (see Del Giudice,, 2021; Binnie et al.,, 2025 for recent reviews) estimator recently proposed in Di Noia et al., (2024). This estimator jointly estimates the ID of the data and the optimal number of NN, ensuring that the underlying probability distribution of the data is approximately uniform in such a neighbourhood. By automatically identifying the optimal locally adaptive neighbourhood structure for each data point, our method adapts to the local structure of the data, enhancing the performance of NN-based projection methods. This adaptive approach can be integrated into several well-known nonlinear projection methods, e.g. LLE, spectral embedding and UMAP, demonstrating significant improvements in their performance. Our method’s adaptability is particularly beneficial in preserving the geometric structure of data, leading to more accurate and meaningful low-dimensional representations.
@@ 2 Methods
In machine learning, most approaches developed in both supervised and unsupervised settings involve several hyper-parameters to be tuned. These can be set a priori by a domain expert based on the available data type, or can be tuned using Hyper-Parameter Optimisation approaches (Yang and Shami,, 2020). Since optimisation algorithms are generally computationally expensive and time-consuming, different techniques have been explored to efficiently search the hyper-parameter space, including Grid Search (LaValle et al.,, 2004), Random Search (Bergstra and Bengio,, 2012), and Bayesian Optimization (Snoek et al.,, 2012). In particular, in dimensionality reduction algorithms, the nature of hyper-parameters is critical, as it significantly influences how observations are mapped into the reduced space and affects the trade-off between preserving the original local and global structure. However, it must be pointed out that hyper-parameter tuning in unsupervised approaches poses a significant challenge, as the absence of ground truth labels prevents direct performance validation (Yang and Shami,, 2020). For instance, in LLE and other dimensionality reduction approaches that leverage a neighbourhood structure for the data, the number of components into which the original observations are mapped (n_components) and the number of neighbours to consider (n_neighbors) are the two main hyper-parameters of the algorithms.
@@ 2.1 Intrinsic dimension and uniform neighbourhoods
Let $X_{1},\dots,X_{n}$ be a random sample taking values in a $D$-dimensional metric space $\mathcal{M}^{D}$. In the following, we use the compact notation $X_{1:n}$ for the sample and $x_{1:n}$ for its realisation. For our purposes, we take $\mathcal{M}^{D}=\mathbb{R}^{D}$ equipped with the Euclidean metric. Moreover, suppose that $X_{1:n}$ are sampled near a locally Euclidean $d$-dimensional manifold embedded in $\mathbb{R}^{D}$ with $d<D$; i.e. the manifold hypothesis is satisfied. Let us consider two open balls $B(x,r_{A})$ and $B(x,r_{B})$ with radii $r_{A}$ and $r_{B}$ such that $r_{A}<r_{B}$ and centred at the same point on the manifold tangent space. Next, we consider a (spatial) Poisson process on $\mathbb{R}^{D}$ with intensity $h:\mathbb{R}^{D}\to\mathbb{R}_{+}$ and we assume that the process is locally homogeneous on $B(x,r_{B})$. It follows that the number of realizations belonging to $B(x,r_{B})$, denoted by $k_{B}$, is Poisson distributed with parameter $\rho\mu(B(x,r_{B}))$, where $\mu$ is the Lebesgue measure on $\mathbb{R}^{D}$ and $\rho$ is the value attained by the intensity function $h$ on the set $B(x,r_{B})$. As a consequence, $k_{A}|k_{B}\sim\mathrm{Binomial}(k_{B},p)$ where $p=(r_{A}/r_{B})^{d}=\tau^{d}$, where $d<D$ is the dimension of the locally Euclidean manifold near which $X_{1:n}$ are sampled. Now, let $k_{A,i}$ and $k_{B,i}$ be, respectively, the number of points in the open balls $B(x_{i},r_{A})$ and $B(x_{i},r_{B})$, where we do not count $x_{i}$ since we condition on it. If $k_{A,1:n}$ are independent conditionally on $k_{B,1:n},\tau$, the likelihood of $d$ reads
$L(d|k_{B,1:n},\tau)=\prod_{i=1}^{n}\binom{k_{B,i}}{k_{A,i}}(\tau^{d})^{k_{A,i}}(1-\tau^{d})^{k_{B,i}-k_{A,i}}$
@@
where $\Omega_{d}=(2\Gamma(3/2))^{d}/\Gamma(d/2+1)$ is the volume of a unit $d$-dimensional hyper-sphere with $\Gamma$ denoting the Euler’s Gamma function. When the underlying Poisson process is homogeneous on $B(x_{i},r_{i,j})$ with intensity constantly equal to $\rho_{i}$, then $v_{i,1},\dots,v_{i,k}$ are independent and identically distributed (i.i.d.) draws from an exponential law with rate $\rho_{i}$. Therefore, the log-likelihood reads
@@
where $q_{1-\alpha,1}$ is the $(1-\alpha)$-quantile of the $\chi_{1}^{2}$ distribution. Note that smaller $\alpha$ implies larger protection against type I error and, consequently, the method tends to identify smaller uniform neighbourhoods. This choice is strongly related to the disentanglement of signal and noise, and it is usually a choice of the researcher. From (2), it is clear that such a selection procedure depends on $d$, however, Di Noia et al., (2024) combined it into an iterative algorithm that starts from an initial estimate of $d$, and improves it through (3) and (1) until convergence. The algorithm is proven to converge at a given tolerance for $n$ large enough, and the resulting ID estimator, named Adaptive BIDE (ABIDE), is proven to be consistent and asymptotically normal. The procedure outputs the estimated ID and the set of optimal $k_{1:n}^{*}$ obtained at the convergence of the iterative procedure, which can be regarded as the identified optimal locally adaptive neighbourhood structure. For the purposes of this paper, unless differently specified, we implicitly consider the ID as the nearest integer to the ABIDE estimate and we denote it by ${d}^{*}$.
@@ 2.2 Adaptive nonparametric dimensionality reduction
As already mentioned, the observed data $X_{1:n}=x_{1:n}$ satisfy the manifold hypothesis. As a consequence, there exists a smooth map $f:\mathbb{R}^{d}\to\mathbb{R}^{D}$ such that $f^{-1}(x_{1:n})=y_{1:n}$ where $y_{1:n}$ is the set of low-dimensional data that we aim to recover. Dimensionality reduction can be seen as the inverse problem of learning the inverse map $g:=f^{-1}$ based on the realisation of a random sample $X_{1:n}$; see Silva and Tenenbaum, (2002). The problem is ill-posed, and it can be solved only by imposing various restrictions on $f$. A huge literature on different approaches naturally stems from this fact. A large number of dimensionality reduction methods are based on performing the inversion, preserving local properties of $f^{-1}$ and of the data. In the realm of local-preserving methods, very successful procedures that have been proposed are based on local neighbourhood structures induced by NN methods. Notable examples are LLE, Spectral Embedding, and Isomap. Such a class of methods can be cast into algorithms of the type
@@ 2.3 Application to LLE: Adaptive LLE
Here, we briefly exemplify how the introduced adaptive framework applies to the LLE method (visually represented in Figure 1). Although several variants have been proposed (e.g. Xue et al.,, 2023; Zhang and Wang,, 2006), we adopt the original formulation of Roweis and Saul, (2000); see Ghojogh et al., (2020) for a review of its variants. The (locally) adaptive LLE method indirectly induces the map $\widehat{g}^{*}$ in (5) through the following algorithm composed of three steps:
@@ 3.1 Datasets
Simulated Manifolds. We consider $n=5100$ points confined on three different manifolds in the 3D space: a torus, a spiral, and a sphere. Subsequently, we generate $D=20$ Gaussian variables. The first three of these variables are added to the manifold points, while the other 17 become additional (noise) coordinates containing. In total, there are 5100 observations with 20 variables, of which we already know the true de-noised ID ($d=3$). This synthetic dataset allows us to test the method’s ability in a controlled environment where we already know the labels and the ID of the manifolds.
@@
Table 1: Main characteristics of the datasets used, namely the data type, number of variables, number of observations, number of labels, the estimated ID $d^{*}$, the standard deviation of the ID estimator, and the ABIDE’s execution time, which represents the additional computational cost introduced by the adaptive mechanism compared to its non-adaptive counterpart.
In Table 1, we also reported $\Delta_{time}$ which represents to the execution time of ABIDE required to estimate the optimal number of neighbours and ID. The value $\Delta_{time}$ therefore represents the additional computational cost, measured on a Mac equipped with an M2 chip, introduced by the adaptive versions with respect to the non-adaptive ones.
@@ 3.2.2 Quantitative Evaluation
As mentioned, the proposed approach allows for autonomously optimising the hyper-parameters by exploiting the locally adaptive neighbourhood structure of the data, independently finding the best hyper-parameters for the number of neighbours (n_neighbors) and the dimension on which to map the original variables (n_components).
@@
In Table 2, we report the results of LLE* together with those obtained using the default hyper-parameter configuration of the standard LLE implementation from the scikit-learn Python package and the modified version of LLE (Zhang and Wang,, 2006) . This comparison is motivated by the practical scenario in which a practitioner, unaware of the most suitable hyper-parameter settings, would typically rely on the default configuration provided by standard software packages.
@@
Another interesting observation is that when we fix the number of neighbours as the mean or the median - or we consider a neighborhood around these summary statistics (represented by the gray area in Figure 3)- of the $k^{*}$ distribution, the results obtained for some metrics are comparable to the results obtained by varying the number of neighbours for the MNIST, Iris, and News Articles datasets. Regarding the Manifolds dataset, this trend is not visible: as previously discussed, looking at Figure 2, we can see how the distribution of $k^{*}$ is much more variable compared to that of the other datasets, which indicates that the data density varies significantly across different regions of the manifold. Therefore, if we take the median of $k^{*}$, we possibly introduce systematic biases due to sub-regions where the density exhibits strong irregularities.
@@
Figure 4: Visualisation of the performance of the 4 considered metrics ARI, Homogeneity, Completeness, and V-Measure for different choices of n_components and n_neighbors in LLE on MNIST dataset. The horizontal dashed lines represent the results of LLE.
For the sake of completeness, in Figure 4, we observe how performance varies as the two hyper-parameters n_components and n_neighbors change for the MNIST dataset. Again, we observe that the choice of the adaptive method leads to better results for almost all the possible configurations of hyper-parameters explored and for all the considered metrics. For some metrics, the best results are obtained with combinations of n_components = 10 and 15 and n_neighbors = 5, values that are comparable to those obtained with the ID estimate ($d^{*}$ = 11.56) and the median number of $k^{*}$ of 7.
@@ 3.2.4 Beyond LLE: Application to UMAP
UMAP (Uniform Manifold Approximation and Projection) is a nonparametric dimensionality reduction method based on manifold learning and topological data analysis (McInnes et al.,, 2018). The algorithm constructs a fuzzy topological representation of the data in the high-dimensional space and then optimizes a low-dimensional embedding to preserve this structure. A critical hyper-parameter in UMAP is the number of nearest neighbors (n_neighbors) used to construct the local neighborhood graph, which directly influences the balance between preserving local and global structure.
We introduce a simple adaptive version of UMAP, denoted as UMAP, that leverages the locally adaptive neighborhood structure $k^{*}_{1:n}$ obtained from the ABIDE estimator. Adopting the notations introduced in Section 2 and following the standard UMAP workflow (McInnes et al.,, 2018), we introduce local scaling parameters $\sigma_{i}$ such that
@@
where $\rho_{i}$ is the distance to the distance to first non-coincident nearest neighbor. Next, the local fuzzy weight vectors are computed as
@@
The sparse weight matrix $W$ defines the fuzzy simplicial set used by the UMAP workflow. Thus, following the standard UMAP procedure, $W$ is symmetrized, and its symmetrized version defines the fuzzy topological representation of the data in the high-dimensional space, where the locally varying neighborhood sizes $k^{*}_{1:n}$ naturally adapt to regions of varying data density. Finally, as in UMAP, UMAP constructs a $d^{*}$-dimensional representation by minimizing the fuzzy set cross-entropy between the high-dimensional and low-dimensional fuzzy representations.
@@ 3.3 Supervised learning
For supervised learning tasks, we must address the challenge of extending our adaptive framework to out-of-sample points not present during the training phase. We develop a local reconstruction approach that maintains the adaptive neighbourhood structure for test data. For each test point, ${x}_{test}$, we temporarily augment the training data including it, estimate its optimal neighborhood size, ${k}^{*}_{test}$ using the intrinsic dimension learned during training, and compute locally adaptive reconstruction weights based on its $k_{test}^{*}$ nearest neighbors. The test point is then projected into the embedding space using a weighted combination of its neighbours’ embeddings. See Algorithm 1 for the detailed procedure.
@@
1: Construct augmented dataset $X_{\text{aug}}=\begin{bmatrix}X_{\text{train}}\\ x_{\text{test}}\end{bmatrix}\in\mathbb{R}^{(n+1)\times D}$, initialize data with ${X}_{\text{aug}}$ Dimension and Neighborhood Estimation:
2: Apply ABIDE: set intrinsic dimension to $d^{*}$ and the optimal neighborhood for ${x}_{\text{test}}$ as the $k^{*}_{\text{test}}$ NN of ${x}_{\text{test}}$ in ${X}_{\text{train}}$ denoted as ${x}_{\text{test}}^{(1)},\ldots,{x}_{\text{test}}^{(k^{*}_{\text{test}})}$Local Linear Reconstruction:

### v2

@@ 1.2 Contributions
In this paper, we introduce a novel nonparametric dimensionality reduction framework that leverages the intrinsic dimension (ID) (see Del Giudice, (2021); Binnie et al., (2025) for recent reviews) estimator recently proposed in Di Noia et al., (2024). This estimator jointly estimates the ID of the data and the optimal number of NN, ensuring that the underlying probability distribution of the data is approximately uniform in such a neighbourhood. By automatically identifying the optimal locally adaptive neighbourhood structure for each data point, our method adapts to the local structure of the data, enhancing the performance of NN-based projection methods. This adaptive approach can be integrated into several well-known nonlinear projection methods, e.g. LLE, spectral embedding and UMAP, demonstrating significant improvements in their performance. Our method’s adaptability is particularly beneficial in preserving the geometric structure of data, leading to more accurate and meaningful low-dimensional representations.
@@ 2 Methods
In machine learning, most approaches developed in both supervised and unsupervised settings involve several hyper-parameters to be tuned. These can be set a priori by a domain expert based on the available data type, or can be tuned using Hyper-Parameter Optimisation approaches (Yang and Shami,, 2020). Since optimisation algorithms are generally computationally expensive and time-consuming, different techniques have been explored to efficiently search the hyper-parameter space, including Grid Search (LaValle et al.,, 2004), Random Search (Bergstra and Bengio,, 2012), and Bayesian Optimization (Snoek et al.,, 2012). In particular, in dimensionality reduction algorithms, the nature of hyper-parameters is critical, as it significantly influences how observations are mapped into the reduced space and affects the trade-off between preserving the original local and global structure. However, it must be pointed out that hyper-parameter tuning in unsupervised approaches poses a significant challenge, as the absence of ground truth labels prevents direct performance validation (Yang and Shami,, 2020). For instance, in LLE and other dimensionality reduction approaches that leverage a neighbourhood structure for the data, the number of components into which the original observations are mapped (n_components) and the number of neighbours to consider (n_neighbours) are the two main hyper-parameters of the algorithms.
@@ 2.1 Intrinsic dimension and uniform neighbourhoods
Let $X_{1},\dots,X_{n}$ be a random sample taking values in $\mathbb{R}^{D}$. In the following, we use the compact notation $X_{1:n}$ for the sample and $x_{1:n}$ for its realisation. Moreover, suppose that $X_{1:n}$ are sampled in a small neighbourhood of a $d$-dimensional $C^{1}$ manifold embedded in $\mathbb{R}^{D}$ with $d<D$; i.e. the manifold hypothesis is satisfied. Let us consider two open balls $B(x,r_{A})$ and $B(x,r_{B})$ with radii $r_{A}$ and $r_{B}$ such that $r_{A}<r_{B}$ and centred at the same point on the manifold tangent space. Next, we consider a (spatial) Poisson process on $\mathbb{R}^{D}$ with intensity $h:\mathbb{R}^{D}\to\mathbb{R}_{+}$ and locally homogeneous on $B(x,r_{B})$. It follows that the number of realizations belonging to $B(x,r_{B})$, denoted by $k_{B}$, is Poisson distributed with parameter $\rho\mu(B(x,r_{B}))$, where $\mu$ is the Lebesgue measure on $\mathbb{R}^{D}$ and $\rho$ is the value attained by the intensity function $h$ on the set $B(x,r_{B})$. As a consequence, $k_{A}|k_{B}\sim\mathrm{Binomial}(k_{B},p)$, where, for a sufficiently small outer ball $B$, $p\approx(r_{A}/r_{B})^{d}=\tau^{d}$, with $d$ corresponding to the ID. Now, let $k_{A,i}$ and $k_{B,i}$ be, respectively, the number of points in the open balls $B(x_{i},r_{A})$ and $B(x_{i},r_{B})$, where we do not count $x_{i}$ since we condition on it. If $k_{A,1:n}$ are independent conditionally on $k_{B,1:n},\tau$, the likelihood of $d$ reads
$L(d)=\prod_{i=1}^{n}\binom{k_{B,i}}{k_{A,i}}(\tau^{d})^{k_{A,i}}(1-\tau^{d})^{k_{B,i}-k_{A,i}}$
@@
where $\Omega_{d}=(2\Gamma(3/2))^{d}/\Gamma(d/2+1)$ is the volume of a unit $d$-dimensional hyper-sphere with $\Gamma$ denoting the Euler’s Gamma function. When the underlying Poisson process is homogeneous on $B(x_{i},r_{i,j})$ with intensity constantly equal to $\rho_{i}$, we model $v_{i,1},\dots,v_{i,k}$ as independent and identically distributed draws from an exponential law with rate $\rho_{i}$. Therefore, the log-likelihood reads
@@
where $q_{1-\alpha,1}$ is the $(1-\alpha)$-quantile of the $\chi_{1}^{2}$ distribution. Note that small values of $\alpha$ imply the identification of larger neighbourhoods. This choice is strongly related to the disentanglement of signal and noise, and it is usually a choice of the researcher.
From (2), it is clear that such a selection procedure depends on $d$, however, Di Noia et al., (2024) combined it into an iterative algorithm that starts from an initial estimate of $d$, and improves it through (3) and (1) until convergence. Under mild conditions, the algorithm is shown to terminate with small numerical tolerance for $n$ large, and the resulting ID estimator is named Adaptive BIDE (ABIDE). In Di Noia et al., (2024) it is proven to be consistent and an expression for its approximate asymptotic variance is provided. The ABIDE workflow outputs the estimated ID and the set of optimal $k_{1:n}^{*}$ obtained at the convergence of the iterative procedure, which can be regarded as the identified optimal locally adaptive neighbourhood structure. For the purposes of this paper, unless differently specified, we implicitly consider the ID as the nearest integer to the ABIDE estimate and we denote it by ${d}^{*}$.
@@ 2.2 Adaptive nonparametric dimensionality reduction
As already mentioned, the observed data $X_{1:n}=x_{1:n}$ satisfy the manifold hypothesis. As a consequence, there exists a map $f:\mathbb{R}^{d}\to\mathbb{R}^{D}$ such that, locally, $f^{-1}(x_{1:n})=y_{1:n}$, where $y_{1:n}$ is the set of low-dimensional data that we aim to recover. Dimensionality reduction can be seen as the inverse problem of learning the inverse map $g:=f^{-1}$ based on the realisation of a random sample $X_{1:n}$; see Silva and Tenenbaum, (2002). The problem is ill-posed, and it can be approached by imposing various restrictions on $f$. A huge literature on different approaches naturally stems from this fact. A large number of dimensionality reduction methods are based on performing the inversion, preserving local properties of the data and of $g$. In the realm of local-preserving methods, very successful procedures that have been proposed are based on local neighbourhood structures induced by NN methods. Notable examples are LLE, Spectral Embedding, and Isomap. Such a class of methods can be cast into algorithms of the type
@@ 2.3 Application to LLE: Adaptive LLE
Here, we briefly exemplify how the introduced adaptive framework applies to the LLE method (visually represented in Figure 1). Although several variants of LLE have been proposed (e.g. Xue et al., (2023); Zhang and Wang, (2006)), we adopt the original formulation of Roweis and Saul, (2000); see Ghojogh et al., (2020) for a review of its variants. The (locally) adaptive LLE method indirectly induces the map $\widehat{g}^{*}$ in (5) through the following algorithm composed of three steps:
@@ 3.1 Datasets
Simulated Manifolds. We consider $n=5100$ points confined on three different manifolds in the 3D space: a torus, a spiral, and a sphere. Subsequently, we generate $D=20$ Gaussian variables. The first three of these variables are added to the manifold points, while the other 17 become additional (noise) coordinates. In total, there are 5100 observations with 20 variables, of which we already know the true de-noised ID ($d=3$). This synthetic dataset allows us to test the method’s ability in a controlled environment where we already know the labels and the ID of the manifolds.
@@
Table 1: Main characteristics of the datasets used, namely the data type, number of variables, number of observations, number of labels, the estimated ID $d^{*}$, the approximate standard deviation of the ID estimator, and the ABIDE’s execution time (measured in seconds), which represents the additional computational cost introduced by the adaptive mechanism compared to its non-adaptive counterpart.
In Table 1, we also reported $\Delta_{time}$, which represents the execution time of ABIDE (measured in seconds) required to estimate the optimal number of neighbours and the intrinsic dimension. The value $\Delta_{time}$ therefore represents the additional computational cost, measured on a Mac equipped with an M2 chip, introduced by the adaptive versions relative to the non-adaptive ones.
@@ 3.2.2 Quantitative Evaluation
As mentioned, the proposed approach allows for autonomously optimising the hyper-parameters by exploiting the locally adaptive neighbourhood structure of the data, independently finding the best hyper-parameters for the number of neighbours (n_neighbours) and the dimension on which to map the original variables (n_components).
@@
In Table 2, we report the results of LLE* together with those obtained using the default hyper-parameter configuration of the standard LLE implementation from the scikit-learn Python package and the modified version of LLE (Zhang and Wang,, 2006) . This comparison is motivated by the scenario in which a practitioner, unaware of the most suitable hyper-parameter settings, would typically rely on the default configuration provided by standard software packages.
@@
Another interesting observation is that when we fix the number of neighbours as the mean or the median - or we consider a neighbourhood around these summary statistics (represented by the gray area in Figure 3)- of the $k^{*}$ distribution, the results obtained for some metrics are comparable to those obtained by using $k^{*}$ for the MNIST, Iris, and News Articles datasets. Regarding the Manifolds dataset, this trend is not visible: as previously discussed, looking at Figure 2, we can see how the distribution of $k^{*}$ is much more variable compared to that of the other datasets, which indicates that the data density varies significantly across different regions of the manifold. Therefore, if we take the median of $k^{*}$, we possibly introduce systematic biases due to sub-regions where the density exhibits strong irregularities.
@@
Figure 4: Visualisation of the performance of the 4 considered metrics ARI, Homogeneity, Completeness, and V-Measure for different choices of n_components and n_neighbours in LLE on MNIST dataset. The horizontal dashed lines represent the results of LLE.
For the sake of completeness, in Figure 4, we observe how performance varies as the two hyper-parameters n_components and n_neighbours change for the MNIST dataset. Again, we observe that the choice of the adaptive method leads to better results for almost all the possible configurations of hyper-parameters explored and for all the considered metrics. For some metrics, the best results are obtained with combinations of n_components = 10 and 15 and n_neighbours = 5, values that are comparable to those obtained with the ID estimate ($d^{*}$ = 11.56) and the median number of $k^{*}$ of 7.
@@ 3.2.4 Beyond LLE: Application to UMAP
UMAP (Uniform Manifold Approximation and Projection) is a nonparametric dimensionality reduction method based on manifold learning and topological data analysis (McInnes et al.,, 2018). The algorithm constructs a fuzzy topological representation of the data in the high-dimensional space and then optimizes a low-dimensional embedding to preserve this structure. A critical hyper-parameter in UMAP is the number of nearest neighbours (n_neighbours) used to construct the local neighbourhood graph, which directly influences the balance between preserving local and global structure.
We introduce a simple adaptive version of UMAP, denoted as UMAP, that leverages the locally adaptive neighbourhood structure $k^{*}_{1:n}$ obtained from the ABIDE estimator. Adopting the notations introduced in Section 2 and following the standard UMAP workflow (McInnes et al.,, 2018), we introduce local scaling parameters $\sigma_{i}$ such that
@@
where $\rho_{i}$ is the distance to first non-coincident nearest neighbour. Next, the local fuzzy weight vectors are computed as
@@
The sparse weight matrix $W$ defines the fuzzy simplicial set used by the UMAP workflow. Thus, following the standard UMAP procedure, $W$ is symmetrized, and its symmetrized version defines the fuzzy topological representation of the data in the high-dimensional space, where the locally varying neighbourhood sizes $k^{*}_{1:n}$ naturally adapt to regions of varying data density. Finally, as in UMAP, UMAP constructs a $d^{*}$-dimensional representation by minimizing the fuzzy set cross-entropy between the high-dimensional and low-dimensional fuzzy representations.
@@ 3.3 Supervised learning
For supervised learning tasks, we must address the challenge of extending our adaptive framework to out-of-sample points not present during the training phase. We develop a local reconstruction approach that maintains the adaptive neighbourhood structure for test data. For each test point, ${x}_{test}$, we temporarily augment the training data including it, estimate its optimal neighbourhood size, ${k}^{*}_{test}$ using the intrinsic dimension learned during training, and compute locally adaptive reconstruction weights based on its $k_{test}^{*}$ nearest neighbours. The test point is then projected into the embedding space using a weighted combination of its neighbours’ embeddings. See Algorithm 1 for the detailed procedure.
@@
1: Construct augmented dataset $X_{\text{aug}}=\begin{bmatrix}X_{\text{train}}\\ x_{\text{test}}\end{bmatrix}\in\mathbb{R}^{(n+1)\times D}$, initialize data with ${X}_{\text{aug}}$ Dimension and Neighbourhood Estimation:
2: Apply ABIDE: set intrinsic dimension to $d^{*}$ and the optimal neighbourhood for ${x}_{\text{test}}$ as the $k^{*}_{\text{test}}$ NN of ${x}_{\text{test}}$ in ${X}_{\text{train}}$ denoted as ${x}_{\text{test}}^{(1)},\ldots,{x}_{\text{test}}^{(k^{*}_{\text{test}})}$Local Linear Reconstruction: